OpenStudy (gabyramirezzz):

Algebra 2 word problem?

OpenStudy (gabyramirezzz):

Going into a final exam which counts as two grades, Kendra has test scores of 84,78,64, and 88. What score does Kendra need on the final exam in order to have an average of 80?

Awolflover1 (awolflover1):

hmm lets do this let x=score she needs to get 80 right?

OpenStudy (gabyramirezzz):

yes

OpenStudy (gabyramirezzz):

so would you add all the scores together and divide it by 4?

Awolflover1 (awolflover1):

\[\frac{ 2x+84+78+64+88 }{ 6 } =80 \] right?

OpenStudy (gabyramirezzz):

yes, because the final counts as two test scores

OpenStudy (gabyramirezzz):

i think thats what i got wrong before

Awolflover1 (awolflover1):

\[\frac{ 2x+314 }{ 6} = ? \]

Awolflover1 (awolflover1):

if you add it up you get 314 :P

OpenStudy (gabyramirezzz):

oh i see

OpenStudy (gabyramirezzz):

and then you would just multiply the 6 out and get 314 +2x=480, right?

Awolflover1 (awolflover1):

yes

Awolflover1 (awolflover1):

now \[2x = 480 - 314\]

OpenStudy (gabyramirezzz):

so x=83?

Awolflover1 (awolflover1):

yes!!

OpenStudy (gabyramirezzz):

thank you so much!!

Awolflover1 (awolflover1):

you're welcome :-)